Peyton:

Which value of x from the set {4, 5, 6, 7}, makes this equation true? 4(8 − x) = 8

jeweldiaz130:

6

Peyton:

thanks

jeweldiaz130:

ur welcome

sylvester:

4(8 - x) = 8, 32 - 4x = 8, -4x = -24, So x = 6